Price Elasticity of Demand

Price elasticity of demand formula: ED= Q2 - Q1/(Q2 + Q1)/2 divided by P2 - P1/(P2 + P1)/2  (remember that price elasticity measures how the price change of a good effects the demand for that good)

Solve the following problem: A local grocer charges $4 each for watermelons, and sold 200 that week. The following week she raised the price to $6 each and only sold 75. Provide your answer, and indicate if your number showed elasticity, unitary elasticity, or was inelastic.
Answer #1

Here Q2 = 75, Q1 = 200. P2 = $6 and P1= $4, Now we can calculate the elasticity by doing Q2-Q1/(Q2+Q1)/2 divided by P2-P1 / (P2+P1)/2. Now Q2-Q1 = 75-200 =-125, (Q2+Q1)/2= (75+200)/2 =275/2 =137.5 Similarly P2-P1 =$(6-4)=$2 and (P2+P1)/2= $(6+4)/2 =$5. Therefore Ed = -125/137.5 divided by 2/5 = - 625/275 = -2.27. As it is own price elasticity is negative that is why it is negative. The elasticity value is 2.27 which is more than 1. So we can say it is elastic. The demand for watermelon is price elastic. It means if there is 100% change in price quantity will change by 227%.